### v1

# Dynamic Order Template Prediction for Generative Aspect-Based Sentiment Analysis

Abstract
Aspect-based sentiment analysis (ABSA) assesses sentiments towards specific aspects within texts, resulting in detailed sentiment tuples. Previous ABSA models often use static templates to predict all of the elements in the tuples, and these models often fail to accurately capture dependencies between elements. Multi-view prompting method improves the performance of ABSA by predicting tuples with various templates and then ensembling the results. However, this method suffers from inefficiencies and out-of-distribution errors. In this paper, we propose a Dynamic Order Template (DOT) method for ABSA, which dynamically generates necessary views for each instance based on instance-level entropy. Ensuring the diverse and relevant view generation, our proposed method improves F1-scores on ASQP and ACOS datasets while significantly reducing inference time.
Dynamic Order Template Prediction for Generative Aspect-Based Sentiment Analysis
[…]
1 Introduction
Aspect-based sentiment analysis (ABSA) aims to identify sentiment for the aspects in a given text rather than simply classifying the overall sentiment of the entire text. ABSA research evolves to generate quadruples consisting of four elements: 1) Aspect ($A$), 2) Category ($C$) for the type of $A$, 3) Opinion ($O$) for $A$, and 4) Sentiment ($S$) for $A$. Many recent studies such as T5-paraphrase (Zhang et al., 2021b) tackle this problem using generative models. These approaches usually get review sentences as input and output the span of quadruples as fixed order paraphrased form, such as "$C$ is $S$ because $A$ is $O$(Zhang et al., 2021a). However, this static single-order template cannot express the dependence between elements as in Figure 1 due to the autoregressive nature of transformer (Vaswani et al., 2017). Moreover, the model’s output can heavily depend on the order of generating each element Hu et al. (2022a).
Figure 1: Comparison of three different generative ABSA methods. 1) static single-view (T5-paraphrase), 2) static multi-view (MvP), and 3) dynamic-view prediction (ours).
Multi-view prompting (Gou et al., 2023) (MvP) deals with this issue by constructing order templates as a channel for "viewing" different perspectives in a sentence. As shown in Figure 1, MvP permutes all possible element orders and sorts them based on the dataset-level entropy of the pre-trained model. Using this entropy, MvP samples top-k orders and adds these orders as a prompt template. During the inference time, MvP conducts majority votes on generated sentiment tuples with various templates. Through this ensemble approach, MvP utilizes the intuition of solving problems from different views in human reasoning and decision Stanovich and West (2000), resulting in significantly higher performance. However, we find that this static multi-view approach of MvP has several drawbacks: 1) Inefficient: Even for samples where the answer can be easily found and multiple views are not necessary, this method generates the same number of views, resulting in unnecessary computation that increases the inference time. 2) Prone to Distribution shift: MvP uses the number of views k as a hyperparameter, applying the same k value across all datasets during training and inference. However, since the optimal number of ensemble models varies according to the data distribution, it requires manual adjustment of the k value for each dataset Shahhosseini et al. (2022), which hinders the transferability to other datasets.
To resolve the aforementioned shortcomings, we propose a Dynamic Order Template (DOT) method for ABSA that combines the advantages of both single-view and multi-view approaches. By prioritizing multiple views based on instance-level entropy, DOT aims to generate only the necessary number of views for each instance during inference. For an example that contains only one tuple as in Figure  1, DOT dynamically creates only one order template (i.e. view) that is necessary for predicting the tuple. After generating the views, DOT generates tuples using the order templates that correspond to the views. This phase operates in a multi-view manner, enabling us to retain the benefits of previous multi-view methods.
Extensive experiments on five widely used sentiment quadruple prediction datasets, derived from ASQP Pontiki et al. (2016); Zhang et al. (2022) and ACOS (Cai et al., 2021, 2023), demonstrate that our method show state-of-the arts performance with significantly lower inference time compared to previous multi-view inference work. Moreover, we show that our method is robust to distribution shift compared to previous methods.
2 Method
Figure 2: Overview of our proposed two stage method. We use two T5 models for each stage: one for generating views, the other for generating sentiment tuples.
Our proposed Dynamic Order Template (DOT) method is composed of two stages as in Figure 2. The first stage predicts the number of order templates to be used as a prompt. (Sec 2.1) Using the order templates, second stage predicts the sentiment tuples. We train the model to generate distinct tuples for each order template, enabling efficient aggregation. (Sec 2.2) For both stages, we map sentiment tuples ($A$, $C$, $S$, $O$) to marker tokens $[A]$, $[C]$, $[S]$, and $[O]$ respectively. For the cases of the instances that contain multiple sentiment tuples, we indicate each tuple with the respective tokens and concatenate the targets with [SSEP] tokens.
2.1 Stage 1: Generating the Number of Views
We assume that the number of sentiment tuples present for each instance corresponds to the required number of views. In other words, we consider that one separate view is necessary for predicting each tuple. We define this individual view as the prediction order for each element of the sentiment tuple as shown in Figure 2. This allows each prediction order to correspond one-to-one with a sentiment tuple in second stage. We set the required number of sentiment tuples $K_{i}$ in the $i^{th}$ instance.
To predict the number of tuples $K_{i}$, we begin by examining the entropy of all possible views generated through permutations. Specifically, we calculate entropy of each view $v$ in instance-level with vanilla T5  Hu et al. (2022a) by calculating conditional generation probability as follows:
$\mathcal{E}_{i,v}$ $=p(\boldsymbol{y_{i,v}}|\boldsymbol{x_{i}}),$ (1)
where $E_{i,v}$ denotes the entropy score of $v$ in the permutation set for $i^{th}$ instance. $y_{i,v}$ is a permuted sentiment tuple in $i^{th}$ instance, based on $v$. At this time, we note that existing ABSA methods have generally struggled to extract $O$ (Chebolu et al., 2023). Based on this observation, we consider that $O$ might hinder the calculation of entropy score. Hence, we calculate entropy scores for $v$ without $O$ (i.e. ($A$, $C$, $S$)) in this stage. We provide a detailed analysis on excluding [O] in Appendix B.2. After computing the entropy, we sort the views by the entropy in ascending order and map each element into marker tokens to get the ranked view $P^{(1)}_{i}$. And then we sample top $K_{i}$ views for each sample $X_{i}$ and concatenate these views as an order template.
Using the original input sentence $X_{i}$, we train the T5 model to generate the first-stage target $y^{(1)}_{i}$. The format of $y^{(1)}_{i}$ is as follows:
$y^{(1)}_{i}=\begin{aligned} &P^{(1)}_{i,1}\,\text{[SSEP]}\,P^{(1)}_{i,2}\,\text{[SSEP]}\,\ldots\,P^{(1)}_{i,K_{i}},\end{aligned}$
where $P^{(1)}_{i,K_{i}}$ denotes the order template for the $j^{th}$ order template of $i^{th}$ instance in the first stage. We set the loss function to train the T5 model as in Equation (2), where $\left\lvert B\right\rvert$ denotes the batch size of the model. The scaling factor is omitted for simplicity.
$\mathcal{L}_{1}$ $=-\sum_{i=1}^{{\left\lvert B\right\rvert}}\sum_{t=1}^{T}\log p(\boldsymbol{y^{(1)}_{i,t}}|\boldsymbol{x_{i}},\boldsymbol{y^{(1)}_{i,{<t}}})$ (2)
2.2 Stage 2: Sentiment Tuple Generation
In the second stage, the model predicts the sentiment tuple of given instance using the number of generated views in the first stage. Different from the first stage, we need to generate all elements in sentiment quadruples including $O$ in this stage. Hence, we re-rank all views to pick $K_{i}$ views including $O$ (i.e. ($A$, $C$, $S$, $O$)). Here, we use the same ranking strategy as in the first stage using the entropy and denote ranked view set as $P^{(2)}_{i}$. We sample top $K_{i}$ views from ranked view and these views are mapped into sentiment tuples one by one, making the model to learn different order template should generate different tuples. We then concatenate the sampled views and add them as a prompt $P_{i}$ to original input sentence $X_{i}$. We design the second stage target $y^{(2)}_{i}$ by placing the corresponding element next to each marker token [SSEP] within $P_{i}$ as follows.
$P_{i}=\begin{aligned} &P^{(2)}_{i,1}\,\text{[SSEP]}\,P^{(2)}_{i,2}\,\text{[SSEP]}\,\ldots\,P^{(2)}_{i,K_{i}},\end{aligned}$
$y^{(2)}_{i}=\begin{aligned} &P^{(2)}_{i,1}\,tuple_{1}\,\text{[SSEP]}\,\ldots\,P^{(2)}_{i,K_{i}}\,tuple_{K_{i}},\end{aligned}$
where $P^{(2)}_{i,K_{i}}$ represents the order template for the $j^{th}$ order template of $i^{th}$ instance in the second stage, and $tuple_{j}$ is the sentiment tuple correspond to $P^{(2)}_{i,K_{i}}$. We design the loss function for training another T5 model in second stage as follows.
$\mathcal{L}_{2}$ $=-\sum_{i=1}^{{\left\lvert B\right\rvert}}\sum_{t=1}^{T}\log p(\boldsymbol{y^{(2)}_{i,t}}|\boldsymbol{x_{i}},\boldsymbol{P_{i}},\boldsymbol{y^{(2)}_{i,{<t}}})$ (3)
2.3 Two-stage inference
During inference time, two stages are conducted sequentially. In the first stage, we generate the necessary views $\hat{y^{(1)}}$ based on the predicted number of views $\hat{K}$ (i.e. the number of tuples). Given $\hat{K}$, we sample top $\hat{K}$ orders from ranked order set $\hat{P}$ and construct order template $\hat{P}$. Note that we rank orders in full elements for each instance. Lastly, we directly add $\hat{P}$ to inference sentence and generate targeted sentiment tuples in second stage.
3 Experiment
3.1 Benchmark Datasets
We adopt two widely used ABSA datasets: ASQP and ACOS, where the task is to predict sentiment quadruples. For ASQP task, we use rest15(R15) and rest16(R16) datasets released from Pontiki et al. (2016); Zhang et al. (2022). For ACOS task, we use laptop16(Lap) and rest16(Rest) datasets constructed by Cai et al. (2021); Pontiki et al. (2016). Also, we adopt additional ACOS dataset(MR) from MEMD restaurant data Xu et al. (2023) which uses a different source from the previous datasets.
3.2 Baselines
We compare our method against several strong baselines for ABSA as follows. Paraphrase Zhang et al. (2021a) formulates a novel paraphrase generation process for ABSA with a single fixed order. Seq2Path Mao et al. (2022) generates sentiments tuples as multiple paths of a tree, and automatically selects valid one. DLO Hu et al. (2022b) augments data via the multiple order templates. MvP aggregates sentiment tuples generated from different orders of prompts via ensembling. AugABSA Wang et al. (2023) generates a original text based on augmented sentiment quadruples. Also, we benchmark popular LLMs such as GPT-3.5, LLaMa-3 Team (2024), and Mistral-7b Jiang et al. (2023). Detailed setups for LLMs are described in Appendix D.
3.3 Implementation Details
We utilize the pre-trained T5-base (Raffel et al., 2020) model as the backbone for the first stage. We also use the model trained in the first stage as the backbone for the second stage, allowing us to leverage a tuned initial point for the ABSA dataset to have the regularization effect inspired by Fu et al. (2023). Also, we eliminate irregularities in tuples through stop-word filtering in the second stage. Please see Appendix A for more details.
3.4 Results
Methods ASQP ACOS Time(s)
R15 R16 Lap Rest MR
Paraphrase 46.93 57.93 43.51 61.16 57.38 40.63
Seq2Path - - 42.97 58.41 - -
DLO 48.18 59.79 43.64 59.99 57.07 260.74
MvP 51.04 60.39 43.92 61.54 58.12 2161.81
AugABSA 50.01 60.88 - - - -
GPT 3.5-turbo 34.27 36.71 16.00 37.71 - -
LLaMa3 8b 37.52 47.60 40.07 54.06 - -
Mistral 7b 44.14 51.96 39.02 53.02 - -
DOT (Ours) 51.91 61.24 44.92 59.25 58.25 298.17
Table 1: F1 scores for ABSA on five datasets. The best results are in bold and the second best are underlined. We conduct experiments with 5 different seeds and report the average of the outcomes. Time denotes the averaged inference time.
F1 score
We use F1 score, which is a standard metric for ABSA, to measure the performance of the systems. As demonstrated in Table 1, our proposed method outperforms all baselines and achieves state-of-the-art performance across four datasets for ABSA. Our model shows slightly reduced performance on the Rest dataset, which we attribute to the substantial number of implicit aspects and opinions within this dataset. Additionally, the size of the Rest dataset is relatively small, being less than half that of the Lap or MR datasets, which are derived from the same ACOS dataset.
Inference time
We also measure inference time using T5-base model for all baselines. We check inference time for each dataset, and average them. As in Table 1, we dramatically reduce inference time particularly compared to the multi-view methods such as MvP Gou et al. (2023), by predicting solely the necessary number of views for each sample. On the other hand, in terms of single view inferences Zhang et al. (2021a), we significantly improve the F1 score performance while suppressing the rate of increase in inference time. We also provide more details about computing the inference time in Appendix C.
3.5 Analysis
Ablation study
To further investigate the effectiveness of each component of our framework, we conduct an ablation study and present the average F1 score across the datasets in Table 2. We first unify the two stages into one, directly generating multiple order templates and tuples without including order prompting to validate the effect of stage division. Additionally, we evaluate the results of directly using the generated views from the first stage, omitting the sampling of new order templates. Lastly, we exclude the multi-view approach by training and testing our model using the order template with the lowest entropy for each instance. By observing the gaps between these variants with the original model, we verify the effectiveness of each component of our method.
Model Configuration Average F1
Full Model 53.94
w/o stage division 52.73 (-1.21)
w/o re-sampling 52.47 (-1.47)
w/o multi view 53.31 (-0.63)
Table 2: Ablation study for the proposed method.
Distribution Shift
To examine the effect of distribution shift of each model, we conduct an in-depth experiment on cross-dataset evaluation. We group the datasets into two groups based on their source: SemEval Pontiki et al. (2016) (R15, R16, Rest) and Yelp (MEMD). Then we assess the performance between these groups by training on one group and testing on the other in a zero-shot setting. For the MvP model, we vary the number of views used for ensembling into 3, 9, and 15 (default) to measure the sensitivity of this number in static multi-view methods. Additionally, we evaluate T5-paraphrase which uses a static single order. Table 3 demonstrates that our model significantly outperforms the baselines in cross-dataset evaluation. While T5-paraphrase experiences a smaller performance drop compared to the others, it still lags behind our method. In particular, MvP exhibits significant performance degradation, irrespective of the number of views. From these experiments, we show that our model can effectively find the optimal number of views even for the out-of-domain datasets.
Train SemEval Yelp
Test SemEval Yelp Yelp SemEval
Paraphrase 52.38 38.52(-11.86) 57.38 44.88(-12.50)
MvP3 55.62 34.42(-21.20) 57.27 41.72(-15.55)
MvP9 56.89 35.02(-21.87) 56.98 42.52(-14.46)
MvP15 57.66 35.21(-21.45) 58.12 41.94(-16.18)
DOT 57.47 39.88(-17.59) 58.25 46.97(-11.28)
Table 3: Cross-dataset evaluation results for validating the effect of distribution shift.
Figure 3: Case study for three main types of results. Blue one denotes correct, red one denotes incorrect, and the yellow one denotes irrelevant.
Case Study
We conduct a case study and analyze the properties of the outputs generated by the proposed method. As depicted in Figure 3, we classify the output results into three main cases.
The first case involves sentences that do not require multiple views for accurate prediction. For these sentences, our model succeeds in making efficient predictions using only a single view. We observe that this case is the most common type in our study, significantly contributing to the model’s efficiency.
The second shows an example predicts require fewer views, but the example actually requires more views. Our analysis reveals that such cases frequently occurs with implicit $O$. As shown in Table 1, this suggests that our model’s performance might lag behind other baselines on the ACOS Rest16 dataset, which contains many samples with implicit $A$ and $O$. Additionally, the model struggles with predicting infrequent $C$ in the training set. Incorporating the concept of self-information and defining the necessary number of views based on the ’amount of information in a sample’ could effectively address this issue.
The final case involves cases with multiple sentiment tuples and longer lengths. We explain that errors in this scenario stem from two main reasons. Firstly, longer sentences include extended phrases that modify $A$ or $O$. Including all these modifiers as elements often leads to errors, a common problem across different models that requires an alternative solution. Secondly, errors occur when the number of tuples is incorrectly predicted in the first stage. If the predicted number of tuples is insufficient, some target sentiment tuples might be overlooked. Conversely, overestimation leads to the extraction of irrelevant aspects, as depicted in the figure. However, we optimize the first stage to reduce tuple count errors, which helped mitigate performance drops by minimizing incorrectly generated or overlooked tuples.
4 Conclusion
We propose Dynamic Order Template (DOT) method for aspect-based sentiment analysis, addressing inefficiencies and out-of-distribution errors in static multi-view prompting. By dynamically generating necessary views based on instance-level entropy, DOT efficiently predicts the sentiment tuple in each instance. Our experiments on ASQP and ACOS datasets demonstrate that DOT achieves state-of-the-art F1-scores with reduced inference time, effectively balancing the strengths of single and multi-view approaches for ABSA.
Limitation
Our DOT method is highly efficient and powerful, yet it still has several limitations. DOT method consists of two stages: view generation and tuple generation. We train separate models for each task, and these two models perform inference sequentially. This form is not end-to-end, so it is disadvantageous in terms of training time and memory.
Also, since we directly connect first stage and second stage, if any errors occur, the errors may propagate and magnify as it moves to the subsequent stage. However, by splitting the task of ’predicting the appropriate number of tuples’ into two sub-tasks—’predicting the appropriate number of tuples’ and ’accurately predicting the tuples’—it becomes significantly easier to achieve accurate results in both areas, thereby enhancing overall performance in our work.
Ethics Statement
This study utilizes the various datasets for aspect-based sentiment analysis, which are accessible online. Additionally, we have properly cited all the papers and sources referenced in our paper. We plan to release the pre-trained model and the code for training the proposed system.
Acknowledgement
This research was supported by Institute for Information & Communications Technology Planning & Evaluation (IITP) through the Korea government (MSIT) under Grant No. 2021-0-01341 (Artificial Intelligence Graduate School Program (Chung-Ang University)).
Appendix A Detailed Experimental Setups
We use AdamW optimizer Loshchilov and Hutter (2017) with a learning rate of 1e-4 for training two T5 models. We set the batch size to 16 for training and 24 for inference. We train the first stage model for 30 epochs, and train 40 epochs for the second stage. Additionally, we observe that the label of the datasets (i.e. sentiment tuples) irregularly contains stop words. For example, as in the first example of Figure 4, the inclusion of negations in the opinion terms is inconsistent. Also, as in the second example, element tuples sometimes contain ambiguous and meaningless stop words as an element. As a result, the fine-tuned model sometimes generates sentiment tuples containing stop words irregularly. It can yield critical performance degradation, even though they don’t affect the meaning of the sentiment elements. To resolve the problem from stop words, we filter these stop words using nltk package(Farkiya et al., 2015) for both generated results and dataset labels. We use four RTX 4090 GPUs to train and evaluate all of the models.
Figure 4: Two examples of irregularity of stop words. Note that these examples are the not all of the stop-word problems.
Appendix B Depth Analysis on First Stage
B.1 Accuracy on the Number of Views
We assess the accuracy of predicting the correct number of views in the first stage and present the results in Table 4. We evaluate the output by comparing it to the labeled sentiment tuples using RMSE and accuracy. We carefully implement the first stage baselines to compare our method properly as follows: Random: We find that the number of sentiment tuples in training dataset is mostly in range of 1 to 6. For each inference, we randomly sample one of the 6 numbers and compare it with our first stage result. Majority: We also reveal that about 60 percent of labels consist of single tuple. We construct a baseline that predicts only 1 for the number of tuples, to check whether our model has ability to predict the number of sentiment tuples of a sentence. Classification: We adopt the RoBERTa model Liu et al. (2019) to evaluate the results when treating the prediction of the number of views as a sequence classification task. We set the classes based on the number of sentiment tuples. As shown in 5, the distribution of tuple counts is skewed towards the lower end, with instances containing more than seven tuples being nearly non-existent. Consequently, we limit the categories from 1 to 6 and clip instances with 7 or more tuples to 6. Additionally, to address label imbalance, we employ a weighted loss function, where the weights are set as the inverse of the frequency ratio for each category as in Equation (4). We use same notation as in Section 2.1, and $\mathcal{I}\left(\right)$ denotes indicator function. This approach enables the model to effectively classify even the less represented classes.
$\boldsymbol{W_{c}}$ $=\frac{|D|}{\sum_{D}\mathcal{I}(min(\boldsymbol{|y|,6})=\boldsymbol{c})}\quad(c\in[1,6])$ (4)
$\mathcal{L}_{cls}$ $=-\sum_{i=1}^{|B|}\boldsymbol{W_{k_{i}}}\log p(\boldsymbol{k_{i}}|\boldsymbol{x_{i}})\mathcal{I}\left(\boldsymbol{k_{i}\leq 6}\right)$
1111222233334444555566667777000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.8number of sentiment tuplesfrequency ratioASQP-Rest15ASQP-Rest16ACOS-Lap16ACOS-Rest16
Figure 5: Distribution of the number of sentiment tuples. The sources are from training datasets of each task. We normalize each count by dividing it by the total number of data points. The number of tuples is clipped to 7.
B.2 Effect of Element Exclusions
We analyze the impact of excluding various marker tokens, including the [O] token representing opinions, to determine which token exclusions contribute to performance improvements. Additionally, we experiment with cases where no element exclusion is performed. In this section, we have also included the second stage results to provide a detailed comparison of the overall performance.
Methods First stage Second stage
RMSE Acc. F1 score
Random 2.80 18.89 -
Majority 0.99 63.39 -
Classification 0.83 61.90 -
$DoT_{first}$ 0.54 77.83 54.33
exclude $[C]$ 0.54 77.53 53.91
exclude $[A]$ 0.53 77.77 53.71
exclude $[S]$ 0.54 77.65 53.55
full elements 0.55 78.22 53.94
Table 4: First stage results for each main baseline and exclusion of specific tokens. We report average RMSE loss and accuracy for first stage, and F1 score for second stage.
As in Table 4, our proposed method outperforms the other baselines and nearly predicts the actual distribution of sentiment tuples within a small margin of error. This result justifies the direct use of the output from the first stage in the second stage. Also, the overall performance improved as the difficulty of predicting the excluded marker token’s element increased. The results from the first stage do not appear to be strongly correlated with those of the second stage. This emphasizes the importance of using a well-tuned initial model for the second stage.
Appendix C Computing Inference Time
We compare inference times based on view methods across different dataset sizes. The dataset consisted of randomly sampled test data from laptop16, with 200, 400, 600, and 800 samples. The baselines were set as static single view (T5-paraphrase) and static multi view (MvP), with the number of views for the multi view fixed at 15.
200200200200400400400400600600600600800800800800001,00010001{,}0001 , $0002{,}$  [^0] Dataset SizeInference Time (s)Single-viewDynamic-viewMulti-view
Figure 6: Inference time among dataset size for each model.
Figure 6 shows that we not only dramatically reduce inference time of utilizing multi views, but also reduce the rate of increase in inference time with respect to the number of datasets. On the other hand, in terms of single view, we significantly increase F1 performance while suppressing the increase in inference time and the rate of its increase. These results suggest that the efficiency of our method becomes more pronounced as the dataset size increases.
Appendix D Detailed Setups for LLM Experiments
As in Table 1, we perform the ABSA task using the GPT 3.5 Turbo, LLaMa-3 8B, and Mistral 7B models, and compared the results with our DOT model. For the GPT model, we utilize in-context learning Brown et al. (2020). We randomly sample 10 instances and combine them with instruction format, and add it as a prompt. For the other two open-source LLMs, we employ instruction tuning Wei et al. (2021) with the training dataset for fine-tuning, using the same instructions as in GPT prompts. To ensure stable model training during fine-tuning, we utilize the LoRa Hu et al. (2021). We present the specific prompts and framework in Figure 7.
According to the following sentiment elements definition:
- The aspect term refers to a specific feature, attribute, or aspect of a product or service that a user may express an opinion about, the aspect term might be null for implicit aspect.
- The opinion term refers to the sentiment or attitude expressed by a user towards a particular aspect or feature of a product or service, the aspect term might be null for implicit opinion.
- The aspect category refers to the category that aspect belongs to, and the available categories includes: {dataset specific categories}.
- The sentiment polarity refers to the degree of positivity, negativity or neutrality expressed in the opinion towards a particular aspect or feature of a product or service, and the available polarities inlcudes: positive’, negative and neutral’.
Recognize all sentiment elements with their corresponding aspect terms, aspect categories, opinion terms and sentiment polarity in the following text with the format of [(’aspect term’, aspect category’, sentiment polarity’, opinion term’), ...]:
Figure 7: Instruction format for two LLM frameworks. We utilize in-context learning for ChatGPT inference, and instruction-tuning for LLaMa-3 and Mistral inference respectively.

### v2

@@
 Aspect-based sentiment analysis (ABSA) assesses sentiments towards specific aspects within texts, resulting in detailed sentiment tuples.
-Previous ABSA models often use static templates to predict all of the elements in the tuples, and these models often fail to accurately capture dependencies between elements.
-Multi-view prompting method improves the performance of ABSA by predicting tuples with various templates and then ensembling the results.
+Previous ABSA models often used static templates to predict all the elements in the tuples, and these models often failed to accurately capture dependencies between elements.
+Multi-view prompting method improves the performance of ABSA by predicting tuples with various templates and then assembling the results.
 However, this method suffers from inefficiencies and out-of-distribution errors.
-In this paper, we propose a Dynamic Order Template (DOT) method for ABSA, which dynamically generates necessary views for each instance based on instance-level entropy.
-Ensuring the diverse and relevant view generation, our proposed method improves F1-scores on ASQP and ACOS datasets while significantly reducing inference time.
+In this paper, we propose a Dynamic Order Template (DOT) method for ABSA, which dynamically creates an order template that contains only the necessary views for each instance.
+Ensuring the diverse and relevant view generation, our proposed method improves F1 scores on ASQP and ACOS datasets while significantly reducing inference time.***Our implementation is publicly available at https://github.com/imsongpasimin/DOT
 Dynamic Order Template Prediction for Generative Aspect-Based Sentiment Analysis
-[…]
-[…]
+Yonghyun Jun  and Hwanhee Leethanks: Corresponding author. Department of Artificial Intelligence, Chung-Ang University {zgold5670, hwanheelee}@cau.ac.kr
 1 Introduction
-Aspect-based sentiment analysis (ABSA) aims to identify sentiment for the aspects in a given text rather than simply classifying the overall sentiment of the entire text.
+Aspect-based sentiment analysis (ABSA) aims to identify the sentiment of aspects in a given text rather than simply classifying the overall sentiment of the entire text.
 ABSA research evolves to generate quadruples consisting of four elements: 1) Aspect ($A$), 2) Category ($C$) for the type of $A$, 3) Opinion ($O$) for $A$, and 4) Sentiment ($S$) for $A$.
-Many recent studies such as T5-paraphrase (Zhang et al., 2021b) tackle this problem using generative models.
-These approaches usually get review sentences as input and output the span of quadruples as fixed order paraphrased form, such as "$C$ is $S$ because $A$ is $O$(Zhang et al., 2021a).
-However, this static single-order template cannot express the dependence between elements as in Figure 1 due to the autoregressive nature of transformer (Vaswani et al., 2017).
-Moreover, the model’s output can heavily depend on the order of generating each element Hu et al.
-(2022a).
-Figure 1: Comparison of three different generative ABSA methods. 1) static single-view (T5-paraphrase), 2) static multi-view (MvP), and 3) dynamic-view prediction (ours).
+Many recent studies, such as T5-paraphrase, tackle this problem using generative models (Zhang et al., 2021b).
+These approaches usually get review sentences as input and output the span of quadruples in fixed order form, such as "$C$ is $S$ because $A$ is $O$(Zhang et al., 2021a).
+However, this static single-order template cannot express the dependence between elements as in Figure 1 due to the autoregressive nature of the transformer (Vaswani et al., 2017).
+Moreover, the model output can heavily depend on the order of generation of each element Hu et al.
+(2022).
+Figure 1: Comparison of three different generative ABSA methods. 1) static single-view, 2) static multi-view, and 3) dynamic-view prediction (ours).
 Multi-view prompting (Gou et al., 2023) (MvP) deals with this issue by constructing order templates as a channel for "viewing" different perspectives in a sentence.
-As shown in Figure 1, MvP permutes all possible element orders and sorts them based on the dataset-level entropy of the pre-trained model.
+As shown in Figure 1, MvP permutes all possible element orders and sorts them based on the entropy of the pre-trained model at the dataset level.
 Using this entropy, MvP samples top-k orders and adds these orders as a prompt template.
-During the inference time, MvP conducts majority votes on generated sentiment tuples with various templates.
-Through this ensemble approach, MvP utilizes the intuition of solving problems from different views in human reasoning and decision Stanovich and West (2000), resulting in significantly higher performance.
-However, we find that this static multi-view approach of MvP has several drawbacks: 1) Inefficient: Even for samples where the answer can be easily found and multiple views are not necessary, this method generates the same number of views, resulting in unnecessary computation that increases the inference time.
-2) Prone to Distribution shift: MvP uses the number of views k as a hyperparameter, applying the same k value across all datasets during training and inference.
-However, since the optimal number of ensemble models varies according to the data distribution, it requires manual adjustment of the k value for each dataset Shahhosseini et al.
+During inference time, MvP conducts majority votes on generated sentiment tuples with various templates.
+Through this ensemble approach, MvP uses the intuition of solving problems from multiple views in human reasoning and decision Stanovich and West (2000), resulting in enhanced performance.
+However, we find that this static multi-view approach of MvP has several drawbacks: 1) Inefficiency: Even for samples where the answer can be easily found and multiple views are not necessary, this method generates the same number of views, resulting in unnecessary computation that increases the inference time.
+2) Limited Transferability: MvP uses the number of views k as a hyperparameter, applying the same k value across all datasets during training and inference.
+However, since the optimal number of ensemble models varies according to the data domain, it requires manual adjustment of the k value for each dataset Shahhosseini et al.
 (2022), which hinders the transferability to other datasets.
@@
 By prioritizing multiple views based on instance-level entropy, DOT aims to generate only the necessary number of views for each instance during inference.
-For an example that contains only one tuple as in Figure  1, DOT dynamically creates only one order template (i.e.
-view) that is necessary for predicting the tuple.
-After generating the views, DOT generates tuples using the order templates that correspond to the views.
+For an example that contains only one tuple, as in Figure 1, DOT dynamically creates only one view as an order template necessary to predict the tuple.
+After generating the views, DOT generates tuples using the views inside the order template.
 This phase operates in a multi-view manner, enabling us to retain the benefits of previous multi-view methods.
-Extensive experiments on five widely used sentiment quadruple prediction datasets, derived from ASQP Pontiki et al.
+Extensive experiments on four widely used sentiment quadruple prediction datasets, derived from ASQP Pontiki et al.
 (2016); Zhang et al.
-(2022) and ACOS (Cai et al., 2021, 2023), demonstrate that our method show state-of-the arts performance with significantly lower inference time compared to previous multi-view inference work.
-Moreover, we show that our method is robust to distribution shift compared to previous methods.
+(2022), ACOS (Cai et al., 2021a, 2023), and MEMD-ABSA Cai et al.
+(2023), demonstrate that our method shows state-of-the-art performance with significantly lower inference time compared to the multi-view approach.
+Moreover, we show that our method is robust to domain shift compared to previous methods, resulting in higher transferability.
 2 Method
-Figure 2: Overview of our proposed two stage method. We use two T5 models for each stage: one for generating views, the other for generating sentiment tuples.
-Our proposed Dynamic Order Template (DOT) method is composed of two stages as in Figure 2.
-The first stage predicts the number of order templates to be used as a prompt.
-(Sec 2.1) Using the order templates, second stage predicts the sentiment tuples.
-We train the model to generate distinct tuples for each order template, enabling efficient aggregation.
-(Sec 2.2) For both stages, we map sentiment tuples ($A$, $C$, $S$, $O$) to marker tokens $[A]$, $[C]$, $[S]$, and $[O]$ respectively.
-For the cases of the instances that contain multiple sentiment tuples, we indicate each tuple with the respective tokens and concatenate the targets with [SSEP] tokens.
-2.1 Stage 1: Generating the Number of Views
-We assume that the number of sentiment tuples present for each instance corresponds to the required number of views.
-In other words, we consider that one separate view is necessary for predicting each tuple.
-We define this individual view as the prediction order for each element of the sentiment tuple as shown in Figure 2.
-This allows each prediction order to correspond one-to-one with a sentiment tuple in second stage.
-We set the required number of sentiment tuples $K_{i}$ in the $i^{th}$ instance.
-To predict the number of tuples $K_{i}$, we begin by examining the entropy of all possible views generated through permutations.
-Specifically, we calculate entropy of each view $v$ in instance-level with vanilla T5  Hu et al.
-(2022a) by calculating conditional generation probability as follows:
-$\mathcal{E}_{i,v}$ $=p(\boldsymbol{y_{i,v}}|\boldsymbol{x_{i}}),$ (1)
-where $E_{i,v}$ denotes the entropy score of $v$ in the permutation set for $i^{th}$ instance.
-$y_{i,v}$ is a permuted sentiment tuple in $i^{th}$ instance, based on $v$.
-At this time, we note that existing ABSA methods have generally struggled to extract $O$ (Chebolu et al., 2023).
-Based on this observation, we consider that $O$ might hinder the calculation of entropy score.
-Hence, we calculate entropy scores for $v$ without $O$ (i.e.
-($A$, $C$, $S$)) in this stage.
-We provide a detailed analysis on excluding [O] in Appendix B.2.
-After computing the entropy, we sort the views by the entropy in ascending order and map each element into marker tokens to get the ranked view $P^{(1)}_{i}$.
-And then we sample top $K_{i}$ views for each sample $X_{i}$ and concatenate these views as an order template.
-Using the original input sentence $X_{i}$, we train the T5 model to generate the first-stage target $y^{(1)}_{i}$. The format of $y^{(1)}_{i}$ is as follows:
+Figure 2: Overview of our proposed two stage method.
+We use two T5 models for each stage: one for generating the initial order template, the other for forming the final order template and generating sentiment tuples.
+[…]
+Also, for the instances that contain multiple sentiment tuples, we indicate each tuple with the respective tokens and concatenate the targets with [SSEP] tokens.
+It is important to note that null or missing values in the datasets are encoded as the literal string "null".
+During template construction, our method maps this string directly to the designated marker tokens, eliminating the need for any special handling.
+2.1 Stage 1: Generating Order Template
+We assume that the number of sentiment tuples $K_{i}$ in $i^{th}$ instance present for each instance corresponds to the required number of views.
+Consequently, we are able to divide the complex sentiment tuple generation task into two relatively simpler subtasks: 1) predicting the number of sentiment tuples in the sentence, and 2) generating sentiment tuples with the exact number of first stage predictions.
+Building separate models specialized for each subtask and enabling collaboration between them was instrumental in achieving significant performance improvements.
+In this framework, each view is interpreted as the ordered scaffold for generating a single sentiment tuple, as depicted in Figure 2.
+This allows each prediction order to correspond one-to-one with a sentiment tuple in the second stage.
+We observe that instead of directly using the value of $K_{i}$ as the target, sampling views corresponding to $K_{i}$ as the target for the model to generate leads to more accurate prediction of the number of tuples in the given instance (More details are in Appendix C).
+To establish the view sampling strategy, we start by ranking all possible views generated through permutations through each entropy score, following  Hu et al.
+(2022).
+Specifically, we calculate entropy of each view $v$ in instance-level with vanilla T5 by calculating conditional generation probability as follows:
+$\mathcal{E}_{i,v}$ $=-\sum P(v|{x_{i}})\log P(v|{x_{i}})$ (1)
+Here, $\mathcal{E}_{i,v}$ is the entropy of the total sequence when the $i_{th}$ instance is input into the T5 model and $v$ is the output.
+At this time, we note that actually utilizing only $A$, $C$, $S$ during the first stage notably facilitates the training process in the second stage.
+We provide a detailed analysis on excluding $O$ in Appendix C.2.
+After computing the entropy, we sort the views by the entropy in ascending order to get the ranked set of views $P^{(1)}_{i}$.
+And then we sample the top $K_{i}$ views for each sample and concatenate these views as an order template.
+Using the original $i^{th}$ input sentence, we train the T5 model to generate the first-stage target $y^{(1)}_{i}$ as follows:
 $y^{(1)}_{i}=\begin{aligned} &P^{(1)}_{i,1}\,\text{[SSEP]}\,P^{(1)}_{i,2}\,\text{[SSEP]}\,\ldots\,P^{(1)}_{i,K_{i}},\end{aligned}$
-where $P^{(1)}_{i,K_{i}}$ denotes the order template for the $j^{th}$ order template of $i^{th}$ instance in the first stage.
+where $P^{(1)}_{i,K_{i}}$ denotes $K_{i}^{th}$ view in $P^{(1)}_{i}$.
 We set the loss function to train the T5 model as in Equation (2), where $\left\lvert B\right\rvert$ denotes the batch size of the model.
@@
 2.2 Stage 2: Sentiment Tuple Generation
-In the second stage, the model predicts the sentiment tuple of given instance using the number of generated views in the first stage.
+In the second stage, the model is trained to generate the sentiment tuple of a given instance using the number of sentiment tuples (i.e.
+$K_{i}$).
 Different from the first stage, we need to generate all elements in sentiment quadruples including $O$ in this stage.
@@
 ($A$, $C$, $S$, $O$)).
-Here, we use the same ranking strategy as in the first stage using the entropy and denote ranked view set as $P^{(2)}_{i}$.
-We sample top $K_{i}$ views from ranked view and these views are mapped into sentiment tuples one by one, making the model to learn different order template should generate different tuples.
-We then concatenate the sampled views and add them as a prompt $P_{i}$ to original input sentence $X_{i}$.
-We design the second stage target $y^{(2)}_{i}$ by placing the corresponding element next to each marker token [SSEP] within $P_{i}$ as follows.
-$P_{i}=\begin{aligned} &P^{(2)}_{i,1}\,\text{[SSEP]}\,P^{(2)}_{i,2}\,\text{[SSEP]}\,\ldots\,P^{(2)}_{i,K_{i}},\end{aligned}$
-$y^{(2)}_{i}=\begin{aligned} &P^{(2)}_{i,1}\,tuple_{1}\,\text{[SSEP]}\,\ldots\,P^{(2)}_{i,K_{i}}\,tuple_{K_{i}},\end{aligned}$
-where $P^{(2)}_{i,K_{i}}$ represents the order template for the $j^{th}$ order template of $i^{th}$ instance in the second stage, and $tuple_{j}$ is the sentiment tuple correspond to $P^{(2)}_{i,K_{i}}$.
-We design the loss function for training another T5 model in second stage as follows.
+Here, we adopt the same strategy as in the first stage, using entropy to form a ranked set of views, $P^{(2)}_{i}$.
+We then sample top $K_{i}$ views from $P^{(2)}_{i}$ and add them as an order template prompt $P_{i}$ to original input sentence.
+We design the second stage target $y^{(2)}_{i}$ by aligning each sentiment tuple with an order template, ensuring that the model learns to generate different tuples for different perspectives.
+Also, we place the corresponding elements next to each marker token within $P_{i}$ as follows:
+$y^{(2)}_{i}=\begin{aligned} &P^{(2)}_{i,1}\otimes tuple_{1}\,\text{[SSEP]}\,\ldots\,P^{(2)}_{i,K_{i}}\otimes tuple_{K_{i}},\end{aligned}$
+where $P^{(2)}_{i,K_{i}}$ represents $K_{i}^{th}$ view in $P^{(2)}_{i}$ and $tuple_{K_{i}}$ is the $K_{i}^{th}$ sentiment tuple for given instance.
+$\otimes$ denotes an interleaved combination between marker tokens and elements.
+Detailed examples for both stages are present in Appendix E.
+We design the loss function for training the T5 model in second stage as follows.
 $\mathcal{L}_{2}$ $=-\sum_{i=1}^{{\left\lvert B\right\rvert}}\sum_{t=1}^{T}\log p(\boldsymbol{y^{(2)}_{i,t}}|\boldsymbol{x_{i}},\boldsymbol{P_{i}},\boldsymbol{y^{(2)}_{i,{<t}}})$ (3)
@@
 During inference time, two stages are conducted sequentially.
-In the first stage, we generate the necessary views $\hat{y^{(1)}}$ based on the predicted number of views $\hat{K}$ (i.e.
-the number of tuples).
-Given $\hat{K}$, we sample top $\hat{K}$ orders from ranked order set $\hat{P}$ and construct order template $\hat{P}$.
-Note that we rank orders in full elements for each instance.
-Lastly, we directly add $\hat{P}$ to inference sentence and generate targeted sentiment tuples in second stage.
+In the first stage, the model generates the initial order template, denoted as $\hat{y^{(1)}}$.
+In the second stage, we count the number of generated views from $\hat{y^{(1)}}$ to set $\hat{K}$.
+Using $\hat{K}$, we sample the top $\hat{K}$ views from the newly ranked set of views and constructs the final order template, referred to as $\hat{P}$.
+Finally, $\hat{P}$ is directly appended to the inference sentence, enabling the generation of different sentiment tuples for each view in $\hat{P}$.
+Throughout inference, we employ a constrained-decoding strategy De Cao et al.
+(2020) to ensure that the output at each stage conforms to the required format.
+The overall two-stage process is described in Figure 2.
 3 Experiment
+Methods ASQP ACOS MEMD Avg Time(s)
+R15 R16 Lap Rest M-Rest M-Lap Books Clothing Hotel
+TAS-BERT 34.78 43.71 27.31 33.53 - - - - - - -
+Extract-Classify 36.42 43.77 35.80 44.61 - - - - - - -
+One-ASQP (large) - - 41.56 60.69 - - - - - - -
+Seq2Path - - 42.97 58.41 - - - - - - -
+AugABSA 50.01 60.88 - - - - - - - - -
+SCRAP 49.93 62.48 - - - - - - - - -
+Paraphrase 46.93 57.93 43.51 61.16 57.38 35.07 39.30 43.00 68.79 50.34 40.63
+DLO 48.18 59.79 43.64 59.99 57.07 35.56 42.63 43.35 70.27 51.16 260.74
+MvP 51.04 60.39 43.92 61.54 58.12 35.25 42.57 43.94 69.06 51.76 2161.81
+GPT-4o 40.45 47.29 24.77 46.53 35.11 20.69 30.39 40.27 24.84 34.48 -
+LLaMa-3.1-8b 37.52 47.60 40.07 54.06 38.10 31.16 28.62 32.21 44.62 39.33 -
+Qwen-2.5-7b 29.93 39.34 12.48 33.56 25.63 24.13 17.77 18.09 38.03 26.66 -
+Mistral-7b 44.14 51.96 39.02 53.02 41.28 26.80 26.54 21.81 40.35 38.32 -
+DOT (Ours) 51.91 61.24 44.92 59.25 58.25 39.02 43.02 43.37 69.94 52.28 298.17
+Table 1: F1 scores for ABSA on nine datasets. Best results are in bold, second-best underlined. Results are averaged over five seeds. Time denotes average inference time.
 3.1 Benchmark Datasets
@@
 For ACOS task, we use laptop16(Lap) and rest16(Rest) datasets constructed by Cai et al.
-(2021); Pontiki et al.
+(2021a); Pontiki et al.
 (2016).
-Also, we adopt additional ACOS dataset(MR) from MEMD restaurant data Xu et al.
-(2023) which uses a different source from the previous datasets.
+Also, we adopt additional ACOS benchmarks from MEMD datasets (Restaurant, Laptop, Books, Clothing, Hotel) Xu et al.
+(2023) which use a different source from the previous datasets.
+We refer to the Restaurant and Laptop datasets in MEMD as M-Rest and M-Laptop, respectively, for the sake of clarity.
 3.2 Baselines
-We compare our method against several strong baselines for ABSA as follows.
-Paraphrase Zhang et al.
-(2021a) formulates a novel paraphrase generation process for ABSA with a single fixed order.
+We benchmark our approach against a suite of extraction- and generation-based baselines.
+For a subset of baselines, we confine evaluation to the four datasets—R15, R16, Laptop, and Restaurant—because these methods were originally optimized for ASQP or ACOS tasks.
+Other baselines are assessed on their full applicable dataset range.
+This selective restriction ensures a fair comparison.
+The complete list of baselines is as follows: TAS-BERT Wan et al.
+(2020) jointly extracts and detects sentimental tuples.
+Extract-Classify Cai et al.
+(2021b) divide the task into two stages: extraction and classification.
+One-ASQP (large) Zhou et al.
+(2023) identify the aspect-opinion-sentiment (AOS) triplets simultaneously.
 Seq2Path Mao et al.
-(2022) generates sentiments tuples as multiple paths of a tree, and automatically selects valid one.
-DLO Hu et al.
-(2022b) augments data via the multiple order templates.
-MvP aggregates sentiment tuples generated from different orders of prompts via ensembling.
+(2022) generates sentiment tuples as multiple paths of a tree, and automatically selects a valid one.
 AugABSA Wang et al.
 (2023) generates a original text based on augmented sentiment quadruples.
-Also, we benchmark popular LLMs such as GPT-3.5, LLaMa-3 Team (2024), and Mistral-7b Jiang et al.
+SCRAP Kim et al.
+(2024) optimizes the model to generate extract-then-assign reasonings and the corresponding sentiment quadruplets in sequence.
+Paraphrase Zhang et al.
+(2021a)  formulates a paraphrase generation process for ABSA with a single fixed order.
+DLO Hu et al.
+(2022)  augments data via the multiple order templates.
+MvP aggregates sentiment tuples generated from multiple orders of prompts via ensembling.
+Also, we benchmark popular LLMs such as GPT-4o Hurst et al.
+(2024), LLaMa-3.1-8b Dubey et al.
+(2024), Qwen-2.5-7b Yang et al.
+(2025), and Mistral-7b Jiang et al.
 (2023).
-Detailed setups for LLMs are described in Appendix D.
+Detailed setups for LLMs are in Appendix H.
 3.3 Implementation Details
@@
 (2023).
-Also, we eliminate irregularities in tuples through stop-word filtering in the second stage.
-Please see Appendix A for more details.
+Additionally, we observe that the label of the datasets (i.e.
+sentiment tuples) irregularly contains stop words.
+Thus, we eliminate irregularities in tuples through stop-word filtering in the second stage.
+We provide a detailed analysis and filtering process in Appendix A.
 3.4 Results
-Methods ASQP ACOS Time(s)
-R15 R16 Lap Rest MR
-Paraphrase 46.93 57.93 43.51 61.16 57.38 40.63
-Seq2Path - - 42.97 58.41 - -
-DLO 48.18 59.79 43.64 59.99 57.07 260.74
-MvP 51.04 60.39 43.92 61.54 58.12 2161.81
-AugABSA 50.01 60.88 - - - -
-GPT 3.5-turbo 34.27 36.71 16.00 37.71 - -
-LLaMa3 8b 37.52 47.60 40.07 54.06 - -
-Mistral 7b 44.14 51.96 39.02 53.02 - -
-DOT (Ours) 51.91 61.24 44.92 59.25 58.25 298.17
-Table 1: F1 scores for ABSA on five datasets.
-The best results are in bold and the second best are underlined.
-We conduct experiments with 5 different seeds and report the average of the outcomes.
-Time denotes the averaged inference time.
-F1 score
-We use F1 score, which is a standard metric for ABSA, to measure the performance of the systems.
-As demonstrated in Table 1, our proposed method outperforms all baselines and achieves state-of-the-art performance across four datasets for ABSA.
-Our model shows slightly reduced performance on the Rest dataset, which we attribute to the substantial number of implicit aspects and opinions within this dataset.
-Additionally, the size of the Rest dataset is relatively small, being less than half that of the Lap or MR datasets, which are derived from the same ACOS dataset.
+Performance Comparison
+We use the F1 score, which is a standard metric for ABSA, to measure the performance of the systems.
+Table 1 demonstrates that our DOT framework achieves state-of-the-art results, ranking first on five of the nine benchmarks and second on three others.
+To further elaborate, a significant factor contributing to the enhanced performance is our approach of dividing the complex sentiment quadruple generation task into two subtasks, each handled by dedicated models.
+As a result, our method achieves better performance compared to a single model handling multi-view processing alone.
+Crucially, the effectiveness of this two-stage approach depends on the accuracy of the first task, which we comprehensively presented in Appendix C.
+However, its performance is slightly lower on the Rest and Clothing datasets, which we analyze in Section F.
 Inference time
-We also measure inference time using T5-base model for all baselines.
-We check inference time for each dataset, and average them.
-As in Table 1, we dramatically reduce inference time particularly compared to the multi-view methods such as MvP Gou et al.
-(2023), by predicting solely the necessary number of views for each sample.
-On the other hand, in terms of single view inferences Zhang et al.
-(2021a), we significantly improve the F1 score performance while suppressing the rate of increase in inference time.
-We also provide more details about computing the inference time in Appendix C.
-3.5 Analysis
-Ablation study
-To further investigate the effectiveness of each component of our framework, we conduct an ablation study and present the average F1 score across the datasets in Table 2.
-We first unify the two stages into one, directly generating multiple order templates and tuples without including order prompting to validate the effect of stage division.
-Additionally, we evaluate the results of directly using the generated views from the first stage, omitting the sampling of new order templates.
-Lastly, we exclude the multi-view approach by training and testing our model using the order template with the lowest entropy for each instance.
-By observing the gaps between these variants with the original model, we verify the effectiveness of each component of our method.
-Model Configuration Average F1
-Full Model 53.94
-w/o stage division 52.73 (-1.21)
-w/o re-sampling 52.47 (-1.47)
-w/o multi view 53.31 (-0.63)
-Table 2: Ablation study for the proposed method.
-Distribution Shift
-To examine the effect of distribution shift of each model, we conduct an in-depth experiment on cross-dataset evaluation.
-We group the datasets into two groups based on their source: SemEval Pontiki et al.
-(2016) (R15, R16, Rest) and Yelp (MEMD).
-Then we assess the performance between these groups by training on one group and testing on the other in a zero-shot setting.
-For the MvP model, we vary the number of views used for ensembling into 3, 9, and 15 (default) to measure the sensitivity of this number in static multi-view methods.
-Additionally, we evaluate T5-paraphrase which uses a static single order.
-Table 3 demonstrates that our model significantly outperforms the baselines in cross-dataset evaluation.
-While T5-paraphrase experiences a smaller performance drop compared to the others, it still lags behind our method.
-In particular, MvP exhibits significant performance degradation, irrespective of the number of views.
-From these experiments, we show that our model can effectively find the optimal number of views even for the out-of-domain datasets.
+We also measure inference time using T5-base model for all baselines. We check inference time for each dataset, and average them.
+As in Table 1, we dramatically reduce inference time particularly compared to the multi-view methods such as MvP Gou et al. (2023), by predicting solely the necessary number of views for each sample.
+On the other hand, single view inferences Zhang et al.
+(2021a); Hu et al.
+(2022) and extraction approaches Wan et al.
+(2020); Cai et al.
+(2021b); Zhou et al.
+(2023) can be more memory- and time-efficientZhou et al.
+(2023), they generally exhibit inferior performance compared to generative methods.
+We provide more details on the inference time in Appendix D.
+Transferability
 Train SemEval Yelp
@@
 DOT 57.47 39.88(-17.59) 58.25 46.97(-11.28)
-Table 3: Cross-dataset evaluation results for validating the effect of distribution shift.
-Figure 3: Case study for three main types of results. Blue one denotes correct, red one denotes incorrect, and the yellow one denotes irrelevant.
-Case Study
+Table 2: Cross-dataset evaluation results for validating the effect of domain shift.
+To examine the transferability of each model, we conduct an in-depth experiment on cross-dataset evaluation.
+We group the datasets into two sources: SemEval Pontiki et al.
+(2016) (R15, R16, Rest) and Yelp (M-Rest), and assess performance by training on one group and testing on the other in a zero-shot setting.
+For the MvP model, we vary the number of views (3, 9, and 15) to evaluate sensitivity in static multi-view methods, while T5-paraphrase uses a static single order.
+As shown in Table 2, our model significantly outperforms the baselines in cross-dataset evaluation.
+Although T5-paraphrase suffers a smaller performance drop, it still falls behind our method, and MvP shows notable degradation regardless of view count.
+These results demonstrate that our model effectively identifies optimal views even for out-of-domain datasets.
+Performance Across Model Scales
+Beyond using T5-base as the sole backbone, we extend our experiments to examine how performance varies with model scale, specifically testing T5-small and T5-large.
+In this setup, we employ the same backbone model for both stages.
+Figure 3: Bar chart illustrating F1 performance variations across T5-small, T5-base, and T5-large backbones on R15, R16, Laptop, and Rest benchmarks.
+As shown in Figure 3, performance consistently improves across all benchmarks as the backbone scale increases, supporting the scaling law Kaplan et al.
+(2020) showing larger models yield better accuracy.
+Nonetheless, these gains do not fully compensate for the inefficiency introduced by loading larger models and performing multiple inference steps.
+Considering the trade-off between accuracy and computational cost, T5-base emerges as the most balanced choice.
+Ablation study
+To further investigate the effectiveness of each component of our framework, we conduct an ablation study and present the results in Table 3.
+Firstly, we record the results without stop-word filtering.
+Also, we unify the two stages into one, directly generating multiple order templates and tuples without including order prompting.
+Additionally, we evaluate the results of sampling the views randomly, checking whether the entropy score is valid.
+Lastly, we exclude the multi-view approach by training and testing our model using only the view with the lowest entropy for each instance as the order template.
+We perform an ablation study by excluding one or more of the four components of our method mentioned earlier.
+By observing the gaps between these variants with the original model, we verify the effectiveness of each component of our method.
+Model Configuration Average F1
+Full Model 54.33
+w/o filtering 53.31 (-1.02)
+w/o stage division 52.73 (-1.60)
+w/o entropy score 52.53 (-1.80)
+w/o multi view 52.31 (-2.02)
+w/o stage division, entropy score 50.04 (-4.29)
+w/o filtering, stage division, entropy score 45.80 (-8.53)
+Table 3: Ablation study for the proposed method, which shows the average F1 across ASQP and ACOS.
+4 Conclusion
+We propose Dynamic Order Template (DOT) method for generative ABSA, addressing inefficiencies and out-of-distribution errors.
+Experiments on nine datasets demonstrate that DOT achieves state-of-the-art performance with reduced inference time, effectively balancing the strengths of previous single and multi-view approaches for ABSA.
+Limitation
+Our DOT method is highly efficient and powerful, yet it still has several limitations.
+DOT method consists of two stages: view generation and tuple generation.
+We train separate models for each task, and these two models perform inference sequentially.
+This form is not end-to-end, so it is disadvantageous in terms of training time and memory.
+Also, since we directly connect first stage and second stage, if any errors occur, the errors may propagate and magnify as it moves to the subsequent stage.
+It results in relatively large standard deviation for different seeds as reported in Table 10.
+However, by splitting the task of ’predicting the appropriate number of tuples’ into two sub-tasks—’predicting the appropriate number of tuples’ and ’accurately predicting the tuples’—it becomes significantly easier to achieve accurate results in both areas, thereby enhancing overall performance in our work.
+Finally, we define the number of necessary views as the number of sentiment tuples for simplicity and efficiency.
+A more complex yet refined method for determining the necessary number of views could be further explored in future research.
+Ethics Statement
+This study utilizes the various datasets for aspect-based sentiment analysis, which are accessible online.
+Additionally, we have properly cited all the papers and sources referenced in our paper.
+We plan to release the pre-trained model and the code for training the proposed system.
+Acknowledgement
+This work was supported by the Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) [RS-2021-II211341, Artificial Intelligence Graduate School Program (Chung-Ang University)].
+Appendix A Detailed Experimental Setup
+We use AdamW optimizer Loshchilov and Hutter (2017) with a learning rate of 1e-4 for training two T5 models.
+We set the batch size to 16 for training and 24 for inference.
+We train the first stage model for 30 epochs, and train 40 epochs for the second stage.
+Figure 4: Two examples of irregularity of stop words. Note that these examples are the not all of the stop-word problems.
+We also note that the dataset labels (i.e., the sentiment tuples) sporadically include stop words.
+For example, as in the first example of Figure 4, the inclusion of negations in the opinion terms is inconsistent.
+Also, as in the second example, element tuples sometimes contain ambiguous and meaningless stop words as elements.
+As a result, the fine-tuned model sometimes generates sentiment tuples containing stop words irregularly.
+It can yield critical performance degradation, even though they don’t affect the meaning of the sentiment elements.
+To resolve the problem of stop words, we filter these stop words using nltk package(Farkiya et al., 2015) for both generated results and dataset labels.
+We use four RTX 4090 GPUs to train and evaluate all of the models.
+Appendix B Case Study
+Figure 5: Case study for the three main types of results. Blue one denotes correct, red one denotes incorrect, and the yellow one denotes irrelevant.
 We conduct a case study and analyze the properties of the outputs generated by the proposed method. As depicted in Figure 5, we classify the output results into three main cases.
@@
 We observe that this case is the most common type in our study, significantly contributing to the model’s efficiency.
-The second shows an example predicts require fewer views, but the example actually requires more views.
+The second shows an example that predicts requires fewer views, but the example actually requires more views.
 Our analysis reveals that such cases frequently occurs with implicit $O$.
@@
 If the predicted number of tuples is insufficient, some target sentiment tuples might be overlooked.
-Conversely, overestimation leads to the extraction of irrelevant aspects, as depicted in the figure.
+Conversely, overestimation leads to the extraction of irrelevant aspects, as depicted in the Figure 5.
 However, we optimize the first stage to reduce tuple count errors, which helped mitigate performance drops by minimizing incorrectly generated or overlooked tuples.
-4 Conclusion
-We propose Dynamic Order Template (DOT) method for aspect-based sentiment analysis, addressing inefficiencies and out-of-distribution errors in static multi-view prompting.
-By dynamically generating necessary views based on instance-level entropy, DOT efficiently predicts the sentiment tuple in each instance.
-Our experiments on ASQP and ACOS datasets demonstrate that DOT achieves state-of-the-art F1-scores with reduced inference time, effectively balancing the strengths of single and multi-view approaches for ABSA.
-Limitation
-Our DOT method is highly efficient and powerful, yet it still has several limitations.
-DOT method consists of two stages: view generation and tuple generation.
-We train separate models for each task, and these two models perform inference sequentially.
-This form is not end-to-end, so it is disadvantageous in terms of training time and memory.
-Also, since we directly connect first stage and second stage, if any errors occur, the errors may propagate and magnify as it moves to the subsequent stage.
-However, by splitting the task of ’predicting the appropriate number of tuples’ into two sub-tasks—’predicting the appropriate number of tuples’ and ’accurately predicting the tuples’—it becomes significantly easier to achieve accurate results in both areas, thereby enhancing overall performance in our work.
-Ethics Statement
-This study utilizes the various datasets for aspect-based sentiment analysis, which are accessible online.
-Additionally, we have properly cited all the papers and sources referenced in our paper.
-We plan to release the pre-trained model and the code for training the proposed system.
-Acknowledgement
-This research was supported by Institute for Information & Communications Technology Planning & Evaluation (IITP) through the Korea government (MSIT) under Grant No.
-2021-0-01341 (Artificial Intelligence Graduate School Program (Chung-Ang University)).
-Appendix A Detailed Experimental Setups
-We use AdamW optimizer Loshchilov and Hutter (2017) with a learning rate of 1e-4 for training two T5 models.
-We set the batch size to 16 for training and 24 for inference.
-We train the first stage model for 30 epochs, and train 40 epochs for the second stage.
-Additionally, we observe that the label of the datasets (i.e.
-sentiment tuples) irregularly contains stop words.
-For example, as in the first example of Figure 4, the inclusion of negations in the opinion terms is inconsistent.
-Also, as in the second example, element tuples sometimes contain ambiguous and meaningless stop words as an element.
-As a result, the fine-tuned model sometimes generates sentiment tuples containing stop words irregularly.
-It can yield critical performance degradation, even though they don’t affect the meaning of the sentiment elements.
-To resolve the problem from stop words, we filter these stop words using nltk package(Farkiya et al., 2015) for both generated results and dataset labels.
-We use four RTX 4090 GPUs to train and evaluate all of the models.
-Figure 4: Two examples of irregularity of stop words. Note that these examples are the not all of the stop-word problems.
-Appendix B Depth Analysis on First Stage
-B.1 Accuracy on the Number of Views
-We assess the accuracy of predicting the correct number of views in the first stage and present the results in Table 4.
-We evaluate the output by comparing it to the labeled sentiment tuples using RMSE and accuracy.
-We carefully implement the first stage baselines to compare our method properly as follows: Random: We find that the number of sentiment tuples in training dataset is mostly in range of 1 to 6.
+Appendix C Depth Analysis on First Stage
+C.1 Accuracy on the Number of Views
+We assess the accuracy of predicting the value of $\hat{K}$ and present the results in Table 4.
+We evaluate the output by comparing it to the number of labeled sentiment tuples using RMSE and accuracy.
+We carefully implement the first stage baselines to compare our method properly as follows: Random: We find that the number of sentiment tuples in the training dataset is mostly in the range of 1 to 6.
 For each inference, we randomly sample one of the 6 numbers and compare it with our first stage result.
-Majority: We also reveal that about 60 percent of labels consist of single tuple.
-We construct a baseline that predicts only 1 for the number of tuples, to check whether our model has ability to predict the number of sentiment tuples of a sentence.
+Majority: We also reveal that about 60 percent of labels consist of a single tuple.
+We construct a baseline that predicts only 1 for the number of tuples, to check whether our model has the ability to predict the number of sentiment tuples of a sentence.
 Classification: We adopt the RoBERTa model Liu et al.
@@
 We set the classes based on the number of sentiment tuples.
-As shown in 5, the distribution of tuple counts is skewed towards the lower end, with instances containing more than seven tuples being nearly non-existent.
+As shown in Figure 6, the distribution of tuple counts is skewed towards the lower end, with instances containing more than seven tuples being nearly non-existent.
 Consequently, we limit the categories from 1 to 6 and clip instances with 7 or more tuples to 6.
 Additionally, to address label imbalance, we employ a weighted loss function, where the weights are set as the inverse of the frequency ratio for each category as in Equation (4).
-We use same notation as in Section 2.1, and $\mathcal{I}\left(\right)$ denotes indicator function.
+We use the same notation as in Section 2.1, and $\mathcal{I}\left(\right)$ denotes the indicator function.
 This approach enables the model to effectively classify even the less represented classes.
@@
 $\mathcal{L}_{cls}$ $=-\sum_{i=1}^{|B|}\boldsymbol{W_{k_{i}}}\log p(\boldsymbol{k_{i}}|\boldsymbol{x_{i}})\mathcal{I}\left(\boldsymbol{k_{i}\leq 6}\right)$
-1111222233334444555566667777000.20.20.20.20.40.40.40.40.60.60.60.60.80.80.80.8number of sentiment tuplesfrequency ratioASQP-Rest15ASQP-Rest16ACOS-Lap16ACOS-Rest16
+11122233344455566677700.20.20.20.40.40.40.60.60.60.80.80.8number of sentiment tuplesfrequency ratioASQP-Rest15ASQP-Rest16ACOS-Lap16ACOS-Rest16
 Figure 6: Distribution of the number of sentiment tuples.
@@
 The number of tuples is clipped to 7.
-B.2 Effect of Element Exclusions
+C.2 Effect of Element Exclusions
 We analyze the impact of excluding various marker tokens, including the [O] token representing opinions, to determine which token exclusions contribute to performance improvements.
@@
 As in Table 4, our proposed method outperforms the other baselines and nearly predicts the actual distribution of sentiment tuples within a small margin of error.
-This result justifies the direct use of the output from the first stage in the second stage.
-Also, the overall performance improved as the difficulty of predicting the excluded marker token’s element increased.
-The results from the first stage do not appear to be strongly correlated with those of the second stage.
-This emphasizes the importance of using a well-tuned initial model for the second stage.
-Appendix C Computing Inference Time
+This result justifies the use of the output from the first stage in the second stage.
+The first stage results in Table 4 do not exhibit significant performance differences among various exclusion.
+However, for the second stage results, which serve as the final output of this task, we observe a significant performance difference.
+The performance in the second stage is generally higher when O is omitted because generating O correctly is the most difficult and crucial task in quadruple prediction (Chebolu et al., 2023).
+If O is not trained in the first stage and is reused in the second stage, the model appears to focus more on learning about O compared to other elements, which already have some level of information.
+Appendix D Computing Inference Time
 We compare inference times based on view methods across different dataset sizes.
@@
 The baselines were set as static single view (T5-paraphrase) and static multi view (MvP), with the number of views for the multi view fixed at 15.
-200200200200400400400400600600600600800800800800001,00010001{,}0001 , $0002{,}$  [^0] Dataset SizeInference Time (s)Single-viewDynamic-viewMulti-view
+20020020040040040060060060080080080001,0001{,}0001 , 0002,0002{,}0002 ,  [^0] Dataset SizeInference Time (s)Single-viewDynamic-viewMulti-view
 Figure 7: Inference time among dataset size for each model.
@@
 These results suggest that the efficiency of our method becomes more pronounced as the dataset size increases.
-Appendix D Detailed Setups for LLM Experiments
-As in Table 1, we perform the ABSA task using the GPT 3.5 Turbo, LLaMa-3 8B, and Mistral 7B models, and compared the results with our DOT model.
+Appendix E Input and Target Examples for Each Stage
+In Figure 8, we provide detailed examples for input and output pairs in each stage.
+The input sentences in the dataset are presented in a basic sentence structure, while the labels consist of lists of sentiment tuples.
+To preprocess this data, during the first stage, the original input sentence is kept unchanged, and the target is set as the initial order template, which consisted of a number of views corresponding to the number of sentiment tuples in the label.
+In the second stage, the input is processed by appending the final order template as a prompt to the original input sentence.
+The target is then constructed by adjusting the order of the elements within the sentiment tuples to align with the corresponding views in the order template.
+Figure 8: Examples for input and target from original dataset for both first and second stage.
+Appendix F Analysis on Implicit Term
+In Table 1, DOT suffers from predicting sentiment tuples in Rest and Clothing domains.
+We noted that the ACOS dataset contains a significant number of instances with implicit aspects or implicit opinions.
+Additionally, we discovered that the Rest and Clothing dataset are smaller in scale compared to other ACOS datasets.
+The scale of each dataset and the number of instances containing implicit terms are recorded in Table 5.
+Based on these observations, we hypothesized that the size of the dataset and the distribution of implicit terms contribute to the performance degradation observed in the Rest and Clothing datasets.
+Datasets ASQP ACOS MEMD
+R15 R16 Lap Rest M-Rest M-Laptop Books Clothing Hotel
+total samples 834 1264 2934 1530 3622 2863 2092 1674 2481
+implicit samples 272 446 1826 822 1801 1751 1523 1083 1278
+implicit sample % 32.6 35.3 62.2 53.7 49.7 61.2 72.8 64.7 51.5
+Table 5: The size of each dataset and the number of samples containing implicit terms.
+For ease of comparison, We also provide the percentage of samples with implicit terms relative to the total number of samples.
+It is evident that the implicit term ratio in the ACOS dataset is higher compared to that in the ASQP dataset.
+Methods Rest M-Rest ¼ M-Rest ½ M-Rest full
+Paraphrase 50.06 40.26 47.77 49.09
+DOT 44.84 35.74 47.81 49.49
+Table 6: F1 scores only for samples containing implicit terms. We report the performance in Rest dataset and the performance trends across different dataset scales.
+As shown in Table 6, it is evident that the F1 score for instances containing implicit terms in the Rest dataset is significantly lower compared to using the paraphrase method.
+Additionally, we observed a performance degradation when training on a randomly selected quarter of the M-Rest dataset.
+However, as the amount of training data from the M-Rest dataset increased, the performance on implicit terms improved, eventually surpassing the F1 score of the paraphrase method in the full M-Rest dataset.
+This result demonstrates that the small size of the dataset with a high proportion of implicit terms is the primary cause of the performance degradation in the Rest and Clothing dataset.
+It also suggests that the performance is likely to improve as the dataset size increases.
+Appendix G Additional Analysis
+In this section, we conduct an in-depth analysis of various aspects of our model.
+For a comprehensive evaluation, we use Paraphrase and MvP as baselines, running identical experiments for comparison.
+We assess performance across multiple tasks using several benchmarks, including R15, R16, Lap, Rest, and M-Rest.
+Different Backbone Model
+We conduct the experiment using different encoder-decoder based model, BART Lewis (2019) as backbone model.
+We utilize BART with the same hyperparameters and data processing techniques applied to the T5 model for three methods including ours.
+However, as in Table 7, we observe a noticeable decline in overall F1-scores for all models, primarily due to insufficient hyperparameter tuning compared to T5.
+Nevertheless, as shown in the results, our method still outperforms the baseline models with BART, suggesting that its effectiveness is not highly dependent on the choice of backbone model.
+Methods ASQP ACOS
+R15 R16 Lap Rest M-Rest
+Paraphrase 31.77 38.15 30.98 36.65 35.74
+MvP 33.48 41.01 32.57 40.40 40.30
+DOT 35.98 41.73 33.12 39.61 40.91
+Table 7: F1 score on benchmark datasets using BART as the backbone model.
+Complex Sentences
+As mentioned in Appendix B, processing long and complex contexts is a well-known challenge, and our model performs similarly to others in this regard.
+We define complex sentences as those containing more than three sentiment tuples, exceeding 22 words in length, or having a Flesch-Kincaid Grade Level Solnyshkina et al.
+(2017) of 9 or higher, representing the top 20% for each criterion.
+We sample these complex sentences and evaluate the F1 scores for these samples.
+As shown in the Table 8, performance degradation in complex sentences is a common issue across all models.
+We attribute the larger performance drop in our model compared to MvP to the fact that it uses fewer views, which limits its capacity to thoroughly analyze complex sentences.
+Nevertheless, our model’s final Complex F1 score remains close to that of MvP and surpasses that of the Paraphrase model.
+Methods ASQP ACOS
+R15 R16 Lap Rest M-Rest
+Paraphrase 44.94 55.73 37.15 56.12 54.37
+MvP 46.71 58.00 37.68 56.06 58.23
+DOT 46.93 57.70 38.57 54.16 57.72
+Table 8: F1 scores evaluated on complex samples only.
+Training Complexity
+Our method may appear complex due to the numerous components that require training.
+However, since our method involves simply training the T5 model twice without complex optimization procedures, the overall training time is not significantly longer than that of other models.
+As shown in the Table  9, even though our model uses 30 and 40 epochs for two stages of training—more than the 20 epochs used in MvP—the total training time remains much shorter than that of MvP.
+In terms of memory usage, only two T5 models are allocated in memory, so the memory consumption does not increase exponentially compared to existing models.
+Methods ASQP ACOS
+R15 R16 Lap Rest M-Rest
+Paraphrase 212.83 314.18 652.31 349.79 815.48
+MvP 3883.74 5008.84 11006.07 6169.02 14634.71
+DOT 1161.73 1648.61 3310.63 1814.41 4157.93
+Table 9: Training duration for each benchmark.
+Standard Deviation
+We conduct experiments using five different random seeds and calculate the standard deviation of the outcomes.
+Results are reported in Table 10.
+Our findings indicate that our model exhibits a higher overall standard deviation compared to other baselines.
+This can be attributed to the structure of the method, where an error at one stage is likely to propagate and accumulate.
+However, it is important to note that the absolute value of the standard deviation is not significantly large.
+In fact, the higher variation suggests that the model may possess greater potential to achieve stronger performance.
+Methods ASQP ACOS
+R15 R16 Lap Rest M-Rest
+Paraphrase $\pm$ 0.44 $\pm$ 0.64 $\pm$ 0.26 $\pm$ 0.68 $\pm$ 0.38
+MvP $\pm$ 0.54 $\pm$ 0.29 $\pm$ 0.48 $\pm$ 0.72 $\pm$ 0.48
+DOT $\pm$ 0.74 $\pm$ 0.85 $\pm$ 1.01 $\pm$ 0.76 $\pm$ 0.42
+Table 10: Standard deviation of outcomes in Table 1.
+Appendix H Detailed Setups for LLM Experiments
+As in Table 1, we perform the ABSA task using the GPT-4o, LLaMa-3.1-8B, and Mistral-7B models, compairing the results with our DOT model.
 For the GPT model, we utilize in-context learning Brown et al.
@@
 We randomly sample 10 instances and combine them with instruction format, and add it as a prompt.
-For the other two open-source LLMs, we employ instruction tuning Wei et al.
+For the other three open-source LLMs, we employ instruction tuning Wei et al.
 (2021) with the training dataset for fine-tuning, using the same instructions as in GPT prompts.
@@
 Recognize all sentiment elements with their corresponding aspect terms, aspect categories, opinion terms and sentiment polarity in the following text with the format of [(’aspect term’, aspect category’, sentiment polarity’, opinion term’), ...]:
-Figure 7: Instruction format for two LLM frameworks. We utilize in-context learning for ChatGPT inference, and instruction-tuning for LLaMa-3 and Mistral inference respectively.
+Figure 9: Instruction format for two LLM frameworks. We utilize in-context learning for GPT-3.5-turbo inference, and instruction-tuning for LLaMa-3.1 and Mistral inference respectively.